OpenStudy (anonymous):

5.In the following equation, identify the x-intercepts in the graph y = 3x^2 + 19x – 40

OpenStudy (saifoo.khan):

\[(x+8) (3 x-5)=0 \]

OpenStudy (saifoo.khan):

\[x = -8\]\[x = \frac 53\]

OpenStudy (anonymous):

if you could show work it would really help me because i have no idea how you got that

OpenStudy (saifoo.khan):

lol, what are u doing back at school? jk.

OpenStudy (anonymous):

its not an in-schol class haha its online, very difficult when you have to teach yourself everything

OpenStudy (anonymous):

*school

OpenStudy (saifoo.khan):

3x^2 + 19x – 40 Factor, 3x^2 + 19x -40 = 0 3x^2 +24x - 6x - 40 = 0 Taking commons, 3x(x+8) -5(x+8) = 0 Taking commons, (3x-5)(x+8) = 0

OpenStudy (saifoo.khan):

lol, online school is very easy!!

OpenStudy (anonymous):

thanks so much! not when its the first time youve done it! atleast not for me, im trying to do my best but some of the material is mind boggling to me. glad people like you are out there to help people like me!

OpenStudy (saifoo.khan):

are u in online college or in online school?

OpenStudy (saifoo.khan):

You welcome! :)

OpenStudy (anonymous):

it can be considered either, its an algebra class, but its my only online class

OpenStudy (saifoo.khan):

i do help many online college/school people who are stuck with their math and stuff.. so was just askin'. :)

OpenStudy (anonymous):

well im sure i will be going to you for help in the future if you are offering it!

OpenStudy (saifoo.khan):

Sure, you can take my email as well.. :)

OpenStudy (anonymous):

alrighty thanks

OpenStudy (saifoo.khan):

No problem!